find an equation of the line having given the slope and containing the given point m=3/4(9,-8)
Answer by Alan3354(69443)

find an equation of the line having given the slope and containing the given point m=3/4(9,-8)
---------------
Use y = mx + b when m = slope and b is the y-intercept.
Find b:
-8 = (3/4)*9 + b
b = -59/4
-----------
--> y = (3/4)x - 59/4
Or 3x - 4y = 59